# From Lipschitz Embedding to Lipschitz Equivalence between Dust-like Self-similar Sets

### v1

@@ 5. Lipschitz equivalence via the homogeneity of $F$
There exists $\alpha\in\mathbb{Z}^{+}$ such that for all $1\leq t\leq m$, $n\geq 1$ and $\bm{i}\in I^{\prime}_{t,n}$, $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ contains at least one cell in $\{F_{\bm{j}}:\bm{j}\in J_{n+\alpha}\}$.
@@
We first explain how Proposition 5.1 leads to the logarithmic commensurability. Assume Proposition 5.1 holds and let $i\in I$ be arbitrary.
For $n\geq 1$, let $k_{n}$ be the unique integer such that $i^{k_{n}}\in I^{\prime}_{1,n}$. So $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ (recall (4.3)). By Proposition 5.1, $\widetilde{f}(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})\supset F_{\bm{j}_{n}}$ for some $\bm{j}_{n}\in J_{n+\alpha}$. Again, we can write as in Lemma 4.3 that
(5.1) $\widetilde{f}^{-1}(F_{\bm{j}_{n}})=\bigcup_{t=1}^{m}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$
where $\mathcal{I}_{t}\subset I^{\prime}_{t,n+\alpha}$. Notice that by Lemma 4.7 and our choice of $\bm{j}_{n}$,
$\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\widetilde{f}(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0\;\textrm{ unless }(t^{\prime},\bm{i}^{\prime})=(1,i^{k_{n}}).$
Thus, from Remark 4.5,
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}})\cap(a_{t^{\prime}}O_{t^{\prime}}K_{\bm{i}^{\prime}}+b_{t^{\prime}}))=0\;\textrm{ unless }(t^{\prime},\bm{i}^{\prime})=(1,i^{k_{n}}),$
which implies that
$\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}}))=\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}})\cap(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})).$
Combining this with (5.1), we have
(5.2) $\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n}}))=\mathcal{H}^{s}\Big{(}\bigcup_{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}}(a_{1}O_{1}K_{\bm{i}}+b_{1})\Big{)}.$
Note that $\{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}\}$ can be written as $\{i^{k_{n}}\omega:\omega\in\Gamma_{n}\}$ for some $\Gamma_{n}\subset I^{*}$. By definition, $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ and $a_{1}r_{i}^{k_{n}}r_{\omega}\approx\delta^{n+\alpha}$ for all $\omega\in\Gamma_{n}$. So there exists a uniform integer $q$ such that $\Gamma_{n}\subset\bigcup_{k=1}^{q}I^{k}$. Then by the pigeonhole principle, we can find $n_{1}<n_{2}$ such that $\Gamma_{n_{1}}=\Gamma_{n_{2}}$. Since $\Psi$ is homogeneous, there are integers $z_{1},z_{2}$ such that $J_{n_{1}+\alpha}=J^{z_{1}}$ and $J_{n_{2}+\alpha}=J^{z_{2}}$. Recalling Proposition 4.4 and (5.2), we have
$\widetilde{c}=\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n_{\ell}}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(a_{1}O_{1}K_{i^{k_{n_{\ell}}}\omega}+b_{1})}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{a_{1}^{s}r_{i}^{k_{n_{\ell}}s}\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(K_{\omega})}{\lambda^{z_{1}s}\mathcal{H}^{s}(F)}$
for $\ell=1,2$. Since $\Gamma_{n_{1}}=\Gamma_{n_{2}}$, we have $r_{i}^{k_{n_{1}}-k_{n_{2}}}=\lambda^{z_{1}-z_{2}}$, i.e., $\frac{\log r_{i}}{\log\lambda}\in\mathbb{Q}$ (choosing $n_{1},n_{2}$ with a large difference, we can require that $z_{1}\neq z_{2}$). Since $i$ is arbitrary, this establish Theorem 1.5.
@@
Since $\Psi$ is homogeneous, for each $n$ we can assign a unique integer $\tau(n)$ such that $J_{n}=J^{\tau(n)}$. For $1\leq t\leq m$ and $n\geq 1$, let
(5.3) $p(t,n):=\max\{p\geq 1:\exists\bm{i}\in I^{\prime}_{t,n}\text{ such that $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for all $\bm{j}\in J^{\tau(n)+p}$}\}$
Suppose on the contrary that Proposition 5.1 fails. Then $\max_{1\leq t\leq m}\sup_{n}p(t,n)=\infty$. Without loss of generality, assume that $\sup_{n}p(1,n)=\infty$.
Let $\gamma\in(0,1)$ be a fixed constant (will be specified later). For $k\geq(\#J)^{6}$, by Szemerédi’s theorem, there is a large integer $n_{k}$ such that any subset of $\{1,\ldots,(\#J)^{p(1,n_{k})}\}$ with cardinality $\geq\gamma(\#J)^{p(1,n_{k})}$ contains a $k$-term arithmetic progression. For notational simplicity, in the rest of the proof we write $p_{k}:=p(1,n_{k})$ and enumerate $J^{p_{k}}$ as $\{\eta_{k,t}:1\leq t\leq(\#J)^{p_{k}}\}$ according to the lexicographical order, that is,
$\eta_{k,1}=\underbrace{1\cdots 1}_{p_{k}\text{ terms}},\,\eta_{k,2}=\underbrace{1\cdots 1}_{p_{k}-1\text{ terms}}\!\!\!\!2,\,\ldots,\,\eta_{k,(\#J)^{p_{k}}}=\underbrace{(\#J)\cdots(\#J)}_{p_{k}\text{ terms}}.$
It will be helpful to note that for any $\bm{j}\in J^{*}$ with $|\bm{j}|\leq p_{k}$,
(5.4) $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}=\{z(\#J)^{p_{k}-|\bm{j}|}+1,z(\#J)^{p_{k}-|\bm{j}|}+2,\ldots,(z+1)(\#J)^{p_{k}-|\bm{j}|}\}$
for some integer $z\geq 0$, i.e., it is a (#J)(\#J)-adic interval of level-$(p_{k}-|\bm{j}|)$.
By (5.3), one can find a word $\bm{i}_{k}\in I^{\prime}_{1,n_{k}}$ with $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ for all $\bm{j}\in J^{\tau(n_{k})+p_{k}}$. On the other hand, recalling Lemma 3.3(3), $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\subset F_{\bm{j}_{k}}$ for some $\bm{j}_{k}\in J_{n_{k}}=J^{\tau(n_{k})}$. Since
$F_{\bm{j}_{k}}=\bigcup_{\theta\in J^{p_{k}}}F_{\bm{j}_{k}\theta}=\bigcup_{1\leq t\leq(\#J)^{p_{k}}}F_{\bm{j}_{k}\eta_{k,t}},$
@@
$\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ $=\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap\bigcup_{1\leq t\leq(\#J)^{p_{k}}}F_{\bm{j}_{k}\eta_{k,t}}$
@@
where $T_{k}:=\{1\leq t\leq(\#J)^{p_{k}}:\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap F_{\bm{j}_{k}\eta_{k,t}}\neq\varnothing\}$. Observe that
@@
$\lesssim(\#T_{k})\cdot\delta^{n_{k}s}\cdot\lambda^{p_{k}s}\mathcal{H}^{s}(F)$ (since $\bm{j}_{k}\in J_{n_{k}}$ and $\eta_{k,t}\in J^{p_{k}}$)
and hence $T_{k}\gtrsim\lambda^{-p_{k}s}=(\#J)^{p_{k}}$. In summary, although every cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{p_{k}}\}$ is not entirely contained in $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$, the latter does intersect a large proportion of these cells.
Picking $0<\gamma<1$ properly at the beginning, we have $T_{k}\geq\gamma(\#J)^{p_{k}}$ for all $k$. As a consequence, $T_{k}$ should contain a $k$-term arithmetic progression, say $AP_{k}$. For the purpose of blowing up on a good sub-cell of $F_{\bm{j}_{k}}$, we perform the following truncation: pick $\theta_{k}\in J^{*}$ such that $F_{\bm{j}_{k}\theta_{k}}$ is a largest cell contained entirely in $\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\bm{j}_{k}\eta_{k,t}}$ and set
@@
which is a subprogression of $AP_{k}$. By (5.4), $\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}$ is simply a #J\#J-adic interval of the form
(5.7) $\{z_{k}(\#J)^{p_{k}-|\theta_{k}|}+1,z_{k}(\#J)^{p_{k}-|\theta_{k}|}+2,\ldots,(z_{k}+1)(\#J)^{p_{k}-|\theta_{k}|}\}$
@@
(5.8) $AP^{\prime}_{k}\geq\frac{\#\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{d_{k}}-2=\frac{(\#J)^{p_{k}-|\theta_{k}|}}{d_{k}}-2.$
It is straightforward to obtain the following lower bound estimate of $AP_{k}^{\prime}$.
@@
$\max AP_{k}-\min AP_{k}\geq(\#AP_{k}-1)d_{k}=(k-1)d_{k}\geq(\#J)^{\ell_{k}}>3\cdot(\#J)^{\ell_{k}-2}.$
Therefore, $\min AP_{k},\max AP_{k}]$ must contain a #J\#J-adic interval
$\{z(\#J)^{\ell_{k}-2}+1,\ldots,(z+1)(\#J)^{\ell_{k}-2}\}$
for some $z\geq 0$. By (5.4), this is just $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}$ for some $\bm{j}\in J^{p_{k}-(\ell_{k}-2)}$. In particular, $F_{\bm{j}}\subset\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\eta_{k,t}}$. Recalling our choice of $\theta_{k}$ (above (5.6)), we have $|\theta_{k}|\leq|\bm{j}|=p_{k}-(\ell_{k}-2)$ and hence
$\#AP^{\prime}_{k}$ $\geq\frac{(\#J)^{\ell_{k}-2}-2d_{k}}{d_{k}}$ (by (5.8))
$=(\#J)^{-2}\cdot(\#J)^{\lfloor\frac{\log(k-1)d_{k}}{\log\#J}\rfloor}\cdot d_{k}^{-1}-2$
$\geq(\#J)^{-3}(k-1)d_{k}\cdot d_{k}^{-1}-2$
$>k(\#J)^{-3}-3\geq k(\#J)^{-4},$
where the last inequality holds since $k\geq(\#J)^{6}$. ∎
@@
For (1), it follows from Lemma 5.3 that
$d_{k}$ $\leq\frac{\#\{1\leq t\leq(\#J)^{p_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{\#AP^{\prime}_{k}-1}$ (by (5.6))
$=\frac{(\#J)^{p_{k}-|\theta_{k}|}}{\#AP^{\prime}_{k}-1}$ (by (5.7))
$\leq(\#J)^{p_{k}-|\theta_{k}|-\lfloor\frac{\log k}{\log\#J}\rfloor-c}$
[…]
$\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})\leq\lambda^{|\bm{j}_{k}|+p_{k}-\widetilde{p}_{k}}=\lambda_{\bm{j}_{k}}\cdot\lambda^{|\theta_{k}|+\lfloor\frac{\log k}{\log\#J}\rfloor+c}=\lambda_{\bm{j}_{k}\theta_{k}}\cdot\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor+c},$
@@
$\operatorname{d_{H}}(\psi_{\bm{j}_{k}\theta_{k}}\widetilde{f}(A_{k,1}),F)=\lambda_{\bm{j}_{k}\theta_{k}}^{-1}\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})=\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor+c},$
@@
For (2), recall that $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ does not contain any cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{p_{k}}\}$ and $|\theta_{k}|\leq p_{k}$. So there are $1\leq t\leq m$, $\bm{i}^{\prime}_{k}\in I^{\prime}_{t,n_{k}}$ with $(t,\bm{i}^{\prime}_{k})\neq(1,\bm{i}_{k})$ such that
$\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}}+b_{t})\cap F_{\bm{j}_{k}\theta_{k}}\neq\varnothing.$
Pick  [^1] For general $\delta$ the word $\bm{j}_{k}\theta_{k}$ might not belong to any $J_{n}$. But one can simply pick $\delta=\lambda$ at the beginning to fix this problem. an integer $q_{k}\in\mathbb{Z}^{+}$ so that $\bm{j}_{k}\theta_{k}\in J_{q_{k}}$. In the spirit of Lemma 4.3, there exists a word $\omega_{k}$ such that $\bm{i}^{\prime}_{k}\omega_{k}\in I^{\prime}_{t,q_{k}}$ and $\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\subset F_{\bm{j}_{k}\theta_{k}}$. Note that
(5.10) $\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\approx\delta^{q_{k}s}\approx\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})=\widetilde{c}^{-1}\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}})),$
@@
for some $0<c_{*}<1$ independent of $k$. This proves (2). ∎

### v2

@@ 5. Lipschitz equivalence via the homogeneity of $F$
There exists $\alpha\in\mathbb{Z}^{+}$ such that for all $1\leq t\leq p$, $n\geq 1$ and $\bm{i}\in I^{\prime}_{t,n}$, $\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ contains at least one cell in $\{F_{\bm{j}}:\bm{j}\in J_{n+\alpha}\}$.
@@
Assume Proposition 5.1 holds and let $i\in I$ be arbitrary.
For $n\geq 1$, let $k_{n}$ be the unique integer such that $i^{k_{n}}\in I^{\prime}_{1,n}$. Proposition 5.1 provides a word $\bm{j}_{n}\in J_{n+\alpha}$ with $F_{\bm{j}_{n}}\subset\widetilde{f}(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})$. By Lemma 4.3, we can write
(5.1) $\widetilde{f}^{-1}(F_{\bm{j}_{n}})=\bigcup_{t=1}^{p}\bigcup_{\bm{i}\in\mathcal{I}_{t}}(a_{t}O_{t}K_{\bm{i}}+b_{t}),$
where $\mathcal{I}_{t}\subset I^{\prime}_{t,n+\alpha}$. For $1\leq t\leq p$, if $\mathcal{I}_{t}\not$$\emptyset$ and $\bm{i}\in\mathcal{I}_{t}$, then letting $\eta$ be the prefix of $\bm{i}$ with $\eta\in I^{\prime}_{t,n}$, we have
$\mathcal{H}^{s}(\widetilde{f}(a_{1}O_{1}K_{i^{k_{n}}}+b_{1})\,\cap\,$ $\widetilde{f}(a_{t}O_{t}K_{\eta}+b_{t}))$
$\geq\mathcal{H}^{s}(F_{\bm{j}_{n}}\cap\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$
$=\mathcal{H}^{s}(\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t}))$ (by (5.1))
[…]
By Lemma 4.7(2), this is impossible unless $t=1$ and $\eta=i^{k_{n}}$. So (5.1) should be
(5.2) $\widetilde{f}^{-1}(F_{\bm{j}_{n}})=\bigcup_{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}}(a_{1}O_{1}K_{\bm{i}}+b_{1}).$
Note that $\{\bm{i}\in\mathcal{I}_{1}:K_{\bm{i}}\subset K_{i^{k_{n}}}\}$ can be written as $\{i^{k_{n}}\omega:\omega\in\Gamma_{n}\}$ for some $\Gamma_{n}\subset I^{*}$. Since $i^{k_{n}}\in I^{\prime}_{1,n}$ and $\mathcal{I}_{1}\subset I^{\prime}_{1,n+\alpha}$, $a_{1}r_{i}^{k_{n}}\approx\delta^{n}$ and $a_{1}r_{i}^{k_{n}}r_{\omega}\approx\delta^{n+\alpha}$ for all $\omega\in\Gamma_{n}$. So there exists an integer $q$ independent of $n$ such that $\Gamma_{n}\subset\bigcup_{k=1}^{q}I^{k}$ for all $n$. By the pigeonhole principle, we may choose $n_{1}<n_{2}$ such that $\Gamma_{n_{1}}=\Gamma_{n_{2}}$. Since $\Psi$ is homogeneous, there are integers $z_{1},z_{2}$ such that $J_{n_{1}+\alpha}=J^{z_{1}}$ and $J_{n_{2}+\alpha}=J^{z_{2}}$. Recalling Proposition 4.4 and (5.2), we have
$\widetilde{c}=\frac{\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{n_{\ell}}}))}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(a_{1}O_{1}K_{i^{k_{n_{\ell}}}\omega}+b_{1})}{\mathcal{H}^{s}(F_{\bm{j}_{n_{\ell}}})}=\frac{a_{1}^{s}r_{i}^{k_{n_{\ell}}s}\sum_{\omega\in\Gamma_{n_{\ell}}}\mathcal{H}^{s}(K_{\omega})}{\lambda^{z_{\ell}s}\mathcal{H}^{s}(F)}$
for $\ell=1,2$. Since $\Gamma_{n_{1}}=\Gamma_{n_{2}}$, we have $r_{i}^{k_{n_{1}}-k_{n_{2}}}=\lambda^{z_{1}-z_{2}}$, i.e., $\frac{\log r_{i}}{\log\lambda}\in\mathbb{Q}$ (choosing $n_{1},n_{2}$ with a large difference, we can require that $z_{1}\neq z_{2}$). Since $i$ is arbitrary, Theorem 1.5 is proved.
@@
Since $\Psi$ is homogeneous, for each $n$ we can assign a unique integer $\tau(n)$ such that $J_{n}=J^{\tau(n)}$. For $1\leq t\leq p$ and $n\geq 1$, let
(5.3) $q(t,n):=\max\{q\geq 1:\exists\bm{i}\in I^{\prime}_{t,n}\text{ such that $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{t}O_{t}K_{\bm{i}}+b_{t})$ for all $\bm{j}\in J^{\tau(n)+q}$}\}.$
Suppose on the contrary that Proposition 5.1 fails. Then $\max_{1\leq t\leq p}\sup_{n}q(t,n)=\infty$. Without loss of generality, assume that $\sup_{n}q(1,n)=\infty$.
Let $\gamma\in(0,1)$ be a fixed constant (will be specified later). For $k\geq(\#J)^{4}$, by Szemerédi’s theorem, there is a large integer $n_{k}$ such that any subset of $\{1,\ldots,(\#J)^{q(1,n_{k})}\}$ with cardinality $\geq\gamma(\#J)^{q(1,n_{k})}$ contains a $k$-term arithmetic progression. For notational simplicity, in the rest of the proof we write $q_{k}:=q(1,n_{k})$ and enumerate $J^{q_{k}}$ as $\{\eta_{k,t}:1\leq t\leq(\#J)^{q_{k}}\}$ according to the lexicographical order, that is,
$\eta_{k,1}=\underbrace{1\cdots 1}_{q_{k}\text{ terms}},\,\eta_{k,2}=\underbrace{1\cdots 1}_{q_{k}-1\text{ terms}}\!\!\!\!2,\,\ldots,\,\eta_{k,(\#J)^{q_{k}}}=\underbrace{(\#J)\cdots(\#J)}_{q_{k}\text{ terms}}.$
It is helpful to note that for any $\bm{j}\in J^{*}$ with $|\bm{j}|\leq q_{k}$,
(5.4) $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}=\{z(\#J)^{q_{k}-|\bm{j}|}+1,z(\#J)^{q_{k}-|\bm{j}|}+2,\ldots,(z+1)(\#J)^{q_{k}-|\bm{j}|}\}$
for some integer $z\geq 0$, i.e., a (#J)(\#J)-adic interval of level-$(q_{k}-|\bm{j}|)$.
By (5.3), one can find a word $\bm{i}_{k}\in I^{\prime}_{1,n_{k}}$ with
(5.5) $F_{\bm{j}}\nsubseteq\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1}),\quad\forall\bm{j}\in J^{\tau(n_{k})+q_{k}}.$
On the other hand, similarly as in (4.4), one can deduce that $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\subset F_{\bm{j}_{k}}$ for some $\bm{j}_{k}\in J_{n_{k}}=J^{\tau(n_{k})}$. Since
$F_{\bm{j}_{k}}=\bigcup_{\theta\in J^{q_{k}}}F_{\bm{j}_{k}\theta}=\bigcup_{1\leq t\leq(\#J)^{q_{k}}}F_{\bm{j}_{k}\eta_{k,t}},$
@@
$\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ $=\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap\bigcup_{1\leq t\leq(\#J)^{q_{k}}}F_{\bm{j}_{k}\eta_{k,t}}$
@@
where $T_{k}:=\{1\leq t\leq(\#J)^{q_{k}}:\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})\cap F_{\bm{j}_{k}\eta_{k,t}}\neq\varnothing\}$. Observe that
@@
$\lesssim(\#T_{k})\cdot\delta^{n_{k}s}\cdot\lambda^{q_{k}s}\mathcal{H}^{s}(F)$ (since $\bm{j}_{k}\in J_{n_{k}}$ and $\eta_{k,t}\in J^{q_{k}}$)
and hence $T_{k}\gtrsim\lambda^{-q_{k}s}=(\#J)^{q_{k}}$. In summary, although every cell in $\{F_{\bm{j}_{k}\theta}:\theta\in J^{q_{k}}\}$ is not entirely contained in $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$, the latter does intersect a large proportion of these cells.
Picking $0<\gamma<1$ properly at the beginning, we have $T_{k}\geq\gamma(\#J)^{q_{k}}$ for all $k$. As a consequence, $T_{k}$ should contain a $k$-term arithmetic progression, say $AP_{k}$. Pick $\theta_{k}\in J^{*}$ such that $F_{\theta_{k}}$ is a largest cell contained entirely in $\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\eta_{k,t}}$ (so $|\theta_{k}|\leq q_{k}$) and set
@@
which is a subprogression of $AP_{k}$. By (5.4), $\{1\leq t\leq(\#J)^{q_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}$ is simply a #J\#J-adic interval of the form
(5.8) $\{z_{k}(\#J)^{q_{k}-|\theta_{k}|}+1,z_{k}(\#J)^{q_{k}-|\theta_{k}|}+2,\ldots,(z_{k}+1)(\#J)^{q_{k}-|\theta_{k}|}\}$
@@
(5.9) $AP^{\prime}_{k}\geq\frac{\#\{1\leq t\leq(\#J)^{q_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{d_{k}}-1=\frac{(\#J)^{q_{k}-|\theta_{k}|}}{d_{k}}-1.$
@@
$\max AP_{k}-\min AP_{k}=(\#AP_{k}-1)d_{k}=(k-1)d_{k}\geq(\#J)^{\ell_{k}}\geq 2\cdot(\#J)^{\ell_{k}-1}.$
Therefore, the interval $\min AP_{k},\max AP_{k}]$ must contain a #J\#J-adic interval of the form
$\big{\{}z(\#J)^{\ell_{k}-1}+1,\ldots,(z+1)(\#J)^{\ell_{k}-1}\big{\}}$
for some $z\geq 0$. By (5.4), this set corresponds to $\{t:F_{\eta_{k,t}}\subset F_{\bm{j}}\}$ for some $\bm{j}\in J^{q_{k}-(\ell_{k}-1)}$. In particular, $F_{\bm{j}}\subset\bigcup_{t=\min AP_{k}}^{\max AP_{k}}F_{\eta_{k,t}}$. Recalling the choice of $\theta_{k}$ (above (5.7)), we have $|\theta_{k}|\leq|\bm{j}|=q_{k}-(\ell_{k}-1)$ and it follows from (5.9) that
$\#AP^{\prime}_{k}$ $\geq\frac{(\#J)^{\ell_{k}-1}}{d_{k}}-1$
$=(\#J)^{-1}\cdot(\#J)^{\lfloor\frac{\log(k-1)d_{k}}{\log\#J}\rfloor}\cdot d_{k}^{-1}-1$
$\geq(\#J)^{-2}(k-1)d_{k}\cdot d_{k}^{-1}-1$
$>k(\#J)^{-2}-2,$
which is at least $k(\#J)^{-3}$ because $k\geq(\#J)^{4}$. ∎
@@
For (1), write $\widetilde{q}_{k}:=q_{k}-|\theta_{k}|-\lfloor\frac{\log k}{\log\#J}\rfloor+4$. Since $k\geq(\#J)^{4}$, Lemma 5.3 implies
$AP_{k}^{\prime}-1\geq(\#J)^{-1}\cdot\#AP_{k}^{\prime}\geq(\#J)^{-4}k\geq(\#J)^{\lfloor\frac{\log k}{\log\#J}\rfloor-4}$
and hence
$d_{k}$ $\leq\frac{\#\{1\leq t\leq(\#J)^{q_{k}}:F_{\eta_{k,t}}\subset F_{\theta_{k}}\}}{\#AP^{\prime}_{k}-1}$ (by (5.7))
$=\frac{(\#J)^{q_{k}-|\theta_{k}|}}{\#AP^{\prime}_{k}-1}$ (by (5.8))
$\leq(\#J)^{q_{k}-|\theta_{k}|-\lfloor\frac{\log k}{\log\#J}\rfloor+4}=(\#J)^{\tilde{q}_{k}}.$
So $AP_{k}$ should intersect every block of $(\#J)^{\widetilde{q}_{k}}$ consecutive integers in $\min AP_{k},\max AP_{k}]$ (and in particular, every #J\#J-adic interval of level-$(q_{k}-\widetilde{q}_{k})$ in $\min AP_{k},\max AP_{k}]$). Since $AP_{k}\subset T_{k}$, by the definition of $T_{k}$ and (5.4), $\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$ (and hence $\widetilde{f}(A_{k,1})$) should intersect every level-$(|\bm{j}_{k}|+q_{k}-\widetilde{q}_{k})$ cell contained in $F_{\bm{j}_{k}\theta_{k}}$. From this, we deduce that
$\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})\leq\lambda^{|\bm{j}_{k}|+q_{k}-\widetilde{q}_{k}}=\lambda_{\bm{j}_{k}}\cdot\lambda^{|\theta_{k}|+\lfloor\frac{\log k}{\log\#J}\rfloor-4}=\lambda_{\bm{j}_{k}\theta_{k}}\cdot\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor-4},$
@@
$\operatorname{d_{H}}(\psi^{-1}_{\bm{j}_{k}\theta_{k}}\widetilde{f}(A_{k,1}),F)=\lambda_{\bm{j}_{k}\theta_{k}}^{-1}\operatorname{d_{H}}(\widetilde{f}(A_{k,1}),F_{\bm{j}_{k}\theta_{k}})=\lambda^{\lfloor\frac{\log k}{\log\#J}\rfloor-4},$
@@
For (2), note that there are $1\leq t\leq p$, $\bm{i}^{\prime}_{k}\in I^{\prime}_{t,n_{k}}$ with $(t,\bm{i}^{\prime}_{k})\neq(1,\bm{i}_{k})$ such that
$\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}}+b_{t})\cap F_{\bm{j}_{k}\theta_{k}}\neq\varnothing;$
otherwise, $F_{\bm{j}_{k}\theta_{k}}\subset\widetilde{f}(a_{1}O_{1}K_{\bm{i}_{k}}+b_{1})$, which contradicts (5.5) because $|\theta_{k}|\leq q_{k}$. Pick any point $x\in a_{t}O_{t}K_{\bm{i}^{\prime}_{k}}+b_{t}$ with $\widetilde{f}(x)\in F_{\bm{j}_{k}\theta_{k}}$. Similarly as in (4.4), we can find a word $\omega_{k}\in I^{*}$ such that $x\in a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t}$, $r_{\bm{i}^{\prime}_{k}\omega_{k}}\approx\lambda_{\bm{j}_{k}\theta_{k}}$ and $\widetilde{f}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\subset F_{\bm{j}_{k}\theta_{k}}$. Note that
(5.11) $\mathcal{H}^{s}(a_{t}O_{t}K_{\bm{i}^{\prime}_{k}\omega_{k}}+b_{t})\approx\mathcal{H}^{s}(F_{\bm{j}_{k}\theta_{k}})=\widetilde{c}^{-1}\mathcal{H}^{s}(\widetilde{f}^{-1}(F_{\bm{j}_{k}\theta_{k}})),$
@@
for some $0<c_{*}<1$ independent of $k$, which completes the proof. ∎